OpenStudy (anonymous):

Given: F^(x) = 5x - 6 and G^(x) = x - 4 F^-1 =

OpenStudy (loser66):

What does F^(x) mean?

OpenStudy (anonymous):

Given: F(x) = 5x - 6 and G(x) = x - 4 F-1 =

OpenStudy (anonymous):

Does that help or no?

OpenStudy (loser66):

no for F-1

OpenStudy (loser66):

f(-1) or \(f^{-1}\)?

OpenStudy (anonymous):

2nd one

OpenStudy (loser66):

From \(f(x) = 5x-6\) when you see x, replace by \(f^{-1}(x)\), when you see f(x), replace by x. What do get?

OpenStudy (anonymous):

Thats the thing i have no idea... i need help learning to do the problem